Archiving on the University of Richmond Physics Department Computer Cluster

G.P.Gilfoyle, August 30, 1999

The University of Richmond Physics Department maintains an array of computers for the research activities of the faculty and students in the Department. The array consists of Sun workstations and Intel-based PC’s running the Linux operating system. To archive data from the array, one of the machines, psc1, has a digital audio tape (DAT) drive and a compact disk (CD) writer. This document is a short description of the use of these devices.

DAT Drive

The DAT drive is a SCSI device located in the lowest expansion slot on the front, righthand-side of psc1. The software device name is /dev/st0.

1. Obtain a digital data storage (DDS) tape. According to the documentation, the DAT drive can read and write DDS-3 tapes though I have only used DDS-2 tapes.

2. The drive is on the bottom of the three expansion slots on the right-hand-side of the front panel of psc1. An arrow on the tape indicates how it should be inserted into the drive. Insert the tape and gently push it in until it the drive engages and completes the tape insertion.

3. Log onto your user account on psc1 (you must have an account on this machine). To position the tape use the mt program. You can find more information using the man pages (i.e., enter ‘man mt’). Some sample commands are listed below.

mt -f /dev/st0 status  Gives the status of the DAT drive.
mt -f /dev/st0 eod  Positions the tape at the end of valid data.

       mt -f /dev/st0 rewind  Rewinds the tape.

       mt -f /dev/st0 offline

Rewinds and unloads the tape. The button on the front panel of the drive does the same thing.

4. Use the tar command to write to the tape. Some sample commands are listed below.

tar -tvf /dev/st0  Lists the files on the tape.

      tar -cvf /dev/st0 courses/

Writes all the data in the ‘courses’ directory to the tape.

      tar -xvf /dev/st0 $PATH/filename

Extracts $PATH/filename from the tape.

CD-Writer

The CD-writer is an IDE device located in the top expansion slot on the front, right-hand-side of psc1. The software device name is /mnt/cdrom.

1. Obtain a rewritable compact disk (CD-RW) or or a recordable one (CD-R). The latter can presumably be written on only once.

2. The drive is on the top of the three expansion slots on the right-hand-side of the front panel of psc1. A button on the right side moves the CD tray in and out. Put the CD on the tray, hit the button, and the tray will move into the drive.

3. Log onto the root account. Note: only the superuser can now burn CD’s. In the future, (i.e. if I ever get around to it), I will set things up so any user can burn CD’s.

4. The CD writing software uses Linux modules that work only if they can avoid conflicts with other modules that address the CD. First make sure that the appropriate module (named ide-cd) has been removed from the kernel by executing the following command.

modprobe -r ide-cd

Note the need for the ‘-r’ option to remove this module from the kernel.

5. Install the kernel module needed for CD writing with the following command.

modprobe ide-scsi

6. Now make an image of the data you want to write to the CD. This step is necessary because once the burning process starts it can’t be stopped without corrupting the CD.

mkisofs -r -o my image directory being archived

7. Write the image file to the CD.

cdrecord -v speed=2 dev=1,0,0 -data my image

This usually takes awhile, but cdrecord prints out a running total of the amount of data that has been written. I’ve also tried speed=4 and it worked fine.

8. Remove the CD-writing module.

modprobe -r ide-scsi

You will now be able to read the CD in the usual manner. Even though you removed the module for reading CD’s (ide-cd, see Step 4), the ide-cd module will be reloaded into the kernel when you mount the cdrom.

Useful Commands for CD-Writing

To see what modules have been installed in the kernel use the lsmod command whose output looks like the following

Module                  Size  Used by
ide-scsi                7072   0
eepro100               12272   2  (autoclean)

where the ide-scsi module has been installed for cd writing. The eepro100 module is always there on psc1 because it is the driver for the ethernet controller.

The command

cdrecord -scanbus

shows the peripheral devices that can use the SCSI protocol. This will include any SCSI devices themselves or a IDE CD writer that uses SCSI emulation. If the ide-scsi module is installed in the kernel (see Step 5 in the previous section), then the output is the following.

Cdrecord release 1.8a24 Copyright (C) 1995-1999 Jrg Schilling
scsibus0:
                  0) 'SEAGATE ' 'ST118202LW      ' '0005' Disk
                  1) 'SEAGATE ' 'ST118202LW      ' '0005' Disk
                  2) 'SEAGATE ' 'DAT    04106-XXX' '7270' Removable Tape
                  3) *
                  4) *
                  5) *
                  6) *
                  7) *
scsibus1:
                100) 'Memorex ' 'CRW-1622        ' 'D3.9' Removable CD-ROM
                101) *
                102) *
                103) *
                104) *
                105) *
                106) *
                107) *

If the ide-scsi module is NOT installed in the kernel the output is only for the devices on scsibus0 and nothing is printed out for scsibus1.

When Disaster Strikes. Currently, the default kernel on psc1 is NOT the one that has cd-writing enabled. This is because of the problems with the automounter that I hope to fix in the future. To use the cd-writing kernel reboot psc1 and at the Lilo prompt enter

LILO: linux1

and follow the instructions above. To get back to the default kernel, just reboot and don’t enter anything.